OpenStudy (anonymous):

Could somebody please help me show my work? I have the right answer the question, I just need to show how I got that answer.. 1. Which values of a and b are a solution to the inequality I 5 - 2a I - B < = 4? Answer: a = 3, b = -1

OpenStudy (vishweshshrimali5):

\(\large{|5-2a| - b \le 4}\) Can you first tell me how you got the answer ?

OpenStudy (vishweshshrimali5):

@Mclovin21 are you there?

OpenStudy (anonymous):

yes, the teacher give us the answers and we have to show the work

OpenStudy (vishweshshrimali5):

Ohh so did you try the question first ?

OpenStudy (anonymous):

yes

OpenStudy (vishweshshrimali5):

Can you show me you work please ?

OpenStudy (anonymous):

I 5 -2a I - b <= 4 5 - 2 x 3- -1 <= 4 5 - 6 - -1 <= 4 -1 - -1 <= 4 0 <= 4 I think it's correct ... but, i'm not sure

OpenStudy (vishweshshrimali5):

@Mclovin21 were there options given in this question ?

OpenStudy (anonymous):

a. a = 6, b = -2 b. a = -4, b = 3 c. a = 3, b = -1 d. a = -3, b = 5

OpenStudy (vishweshshrimali5):

By the way you did one mistake in your solution. \(\color{blue}{\text{Originally Posted by}}\) @Mclovin21 I 5 -2a I - b <= 4 \(\color{red}|5 - 2 * 3\color{red}|-\color{red}( -1\color{red}) \le 4\) 5 - 6 - -1 <= 4 -1 - -1 <= 4 0 <= 4 I think it's correct ... but, i'm not sure \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

so I was correct just a few sign errors

OpenStudy (vishweshshrimali5):

Okay see the symbols I have added in red colour are very important: (1) You forgot to take the absolute value of (5-2a). This made your solution wrong.

OpenStudy (anonymous):

ohhh how do you find the absolute value of it?

OpenStudy (vishweshshrimali5):

The correct solution would have been this: \(\color{blue}{\text{Originally Posted by}}\) @Mclovin21 \(|5 -2a| - b \le 4\) \(\color{red}|5 - 2 * 3\color{red}|- \color{red}(-1\color{red}) \le 4\) \(\color{red}|5 - 6\color{red}| - \color{red}(-1\color{red}) \le 4\) \(\color{red}|-1\color{red}| -\color{red}(-1\color{red}) \le 4\) \(\color{red}1 + 1 \le 4\) \(\color{red}2 \le 4\) I think it's correct ... but, i'm not sure \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

ohhhhh okay! I see it now thank you!

OpenStudy (vishweshshrimali5):

Absolute value is just like a machine in which you insert a number and it gives the positive value of the same number in the output. For example, \(|2| = 2\) (Since, 2 is already positive so it didn't affect it and showed the same number) \(|-2| = 2\) (Since, -2 is negative it changed it to 2\) Basically, \(|x| = x, \text{If x is positive}\) and \(|x| = -x, \text{If x is negative}\) So, in the question you had \(|-1|\) , now, -1 is negative so, the second part of definition will follow, that is \(|-1| = \color{red}-\color{blue}(\color{green}{-1}\color{blue}) = \color{pink}1\)

OpenStudy (vishweshshrimali5):

No problem @Mclovin21 . Best of luck Good day